OpenStudy (anonymous):

Probability Question: A box contains 8 tickets bearing numbers 1,2,3,4,5,6,8,10. One ticket is drawn and kept aside. Then a second ticket is drawn. What's the probability that both tickets show even numbers?

OpenStudy (wolfe8):

Did you know that the probability of 2 consecutive events is the product of the probabilities of each separate event?

OpenStudy (anonymous):

yes plz reply

OpenStudy (anonymous):

plz tell dear

OpenStudy (wolfe8):

Alright so what is the probability of getting an even ticket in the first pull? Then, since you will only have 7 tickets left(the one pulled is put aside), what is the probability of getting another even ticket? Multiply them together.

OpenStudy (anonymous):

1/8 X 1/7 = 0.125 X 0.1428 = 0.1785

OpenStudy (anonymous):

is it correct?

OpenStudy (wolfe8):

Not quite. Since you need even tickets, how many even tickets do you have?

OpenStudy (anonymous):

5 sir

OpenStudy (anonymous):

4

OpenStudy (wolfe8):

Ok list all the even tickets.

OpenStudy (anonymous):

2,4,6,8,10

OpenStudy (wolfe8):

So that's 5. Probability should be 5/8 for the first one. Second one?

OpenStudy (anonymous):

plz tell

OpenStudy (anonymous):

its mean answer is 5/8 x 4/8 =0.3125

OpenStudy (wolfe8):

Well when you pull out an even ticket you will have one less even ticket and one less total ticket. You were almost there last time but you forgot there was more than 1 even ticket

OpenStudy (anonymous):

plz explain in detail

OpenStudy (wolfe8):

So when you pull out the first ticket, your prob. of getting an even one is 5/8. After that, you keep it aside and will have 1 less. So it should be 4/7

OpenStudy (anonymous):

Is it correct to write= P(AnB)= P(A) X P(B) = 5/8 X 4/7

OpenStudy (wolfe8):

Right. I'm not used to using the notations but I believe so.

OpenStudy (anonymous):

Thanks a lot dear my pleasure to have friend like u

OpenStudy (wolfe8):

You're welcome. The pleasure is mine to help someone learn.

OpenStudy (wolfe8):

Please close the question when you are done though.